OpenStudy (anonymous):

Express the following function, F(x) as a composition of two functions f and g. (will draw function F(x) in a sec!) ***Please explain! Thank you =)

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

Any ideas where to start?

OpenStudy (anonymous):

i have no clue about compositions and stuff :/ they confuse me :(

jimthompson5910 (jim_thompson5910):

Say you have f(x) = x^2 and you have g(x) = x+1 Then f(x) = x^2 f(g(x)) = (g(x))^2 f(g(x)) = (x+1)^2

jimthompson5910 (jim_thompson5910):

Notice how I replaced every 'x' with g(x) in the second line

jimthompson5910 (jim_thompson5910):

Then in the third line, I replaced g(x) with what it really was...which is x+1

OpenStudy (anonymous):

i think i get it :) sooo in this problem, is it asking me to solve for f(x) and g(x) ??

jimthompson5910 (jim_thompson5910):

yes, pretty much

OpenStudy (anonymous):

okay :) what do i do first tho? :)

jimthompson5910 (jim_thompson5910):

let's say g(x) = x^2, what would f(x) have to be to make f(g(x)) like what you see above?

OpenStudy (anonymous):

(x+1)^2 ??